please help me solve these equations:
1)p^2+3p-54
2)n^2-n-12
3)a^2+10a+24
i really need help....thank u

yasmine L

Answer by stanbon(75887)

You have not posted equations.
But I will assume you meant to.
--------------
1)p^2+3p-54 = 0
(p+9)(p-6) = 0
p = -9 or p = 6
==========================
2)n^2-n-12 = 0
(n-4)(n+3) = 0
n = 4 or n = -3
==========================
3)a^2+10a+24 = 0
(a+6)(a+4) = 0
a = -4 or a = -6
==========================
Cheers,
Stan H.